# Modeling Light Signals Using Data from the First Pulsed Neutron Source Program at the DUNE Vertical Drift ColdBox Test Facility at CERN Neutrino Platform

### v1

@@ 6.3 Discussion on Excess at High PE in Data
Here we discuss several possible sources that could produce the high PE excess observed in the data in Fig. 6.
@@
Cosmic background. The excess at high PE region could originate from not well modeled cosmic flux from cosmic runs. To test this hypothesis, we performed a fit to the timing distribution separately for events with number of PE below and above 1200 in Fig. 6. For events with PE 1200, the timing distribution would be flat as one would expect for cosmic background. Instead, a decay time profile similar to the events with PE 1200 is observed (Fig. 9). This provides strong evidence that the excess at high PE is from neutron-related signals not modeled in the simulation.
@@
XA Absolute PDE. In the analysis (Sec. 5.2), a relative PDE calibration was performed to the 3% absolute PDE used in the simulation. However, the actual XA absolute PDE can vary by up to 50% when SiPMs are biased at higher overvoltages with a non-negligible cross-talk probability. The absolute PDE is also affected by the detailed XA configuration. A smaller absolute PDE in simulation would underestimate the number of PE and could partially explain the high PE excess in data. This is further discussed in Sec. 7 as a systematic effect. The precise measurement of absolute PDE for XA modules requires a dedicated laboratory setup and analysis. More details can be found in ref. [VD_XA_PDE].
@@ 7 Systematic Errors
Several important systematic effects are studied using the Fluka simulation. The uncertainty of the electric field in the buffer LAr region affects the scintillation photon yield. A simulation with zero electric field ("0 field") in this volume is evaluated, where the light yield reaches a maximum compared to the nominal electric field of 454 V/cm. The uncertainty of the initial production of scintillation light in buffer LAr and propagation to the active LAr readout volume is evaluated by setting the photon production in the buffer LAr volume to zero ("act only"). The uncertainty of the PNS position relative to the active volume inside the CB is modeled by shifting the PNS in the x (drift) and y (horizontal) directions by 5 cm ("dispx" and "dispy"). The uncertainty on the photon detector efficiency is studied by increasing it to 4.44% ("heff"), a value that corresponds to the product of efficiency and cross-talk probability for one of the device configurations described in ref. [VD_XA_PDE].
@@
The peak ADC amplitude to photoelectron calibration constant used in the data discussed in Sec. 5.1 affects the observed amount of light on each XA. The systematic shift of this constant up and down by one ADC unit is applied to the calculation in data, and the resulting counts are presented in Tab. 5. Overall, we observe that the non-cosmic component in data has a max shift of 8.3% to the nominal scenario from C4.
| XA | Cosmics | Inactive LAr (MC) | Active LAr Inelastic (MC) | Active LAr Capture (MC) | Total Predicted (MC) | Data | Data - Cosmics |
| C1 | 106951 | 35511 | 2020 | 24022 | 61552 | 153597 | 46646 |
| C2 | 91296 | 7556 | 238 | 4778 | 12572 | 108528 | 17232 |
| C3 | 88927 | 8381 | 287 | 6599 | 15267 | 104194 | 15267 |
| C4 | 79768 | 137749 | 9347 | 44979 | 192076 | 192214 | 112446 |
| XA | Cosmics | Inactive LAr (MC) | Active LAr Inelastic (MC) | Active LAr Capture (MC) | Total Predicted (MC) | Data | Data - Cosmics |
| C1 | 117967 | 38898 | 2213 | 26313 | 67424 | 171432 | 53465 |
| C2 | 96981 | 8276 | 261 | 5234 | 13771 | 116530 | 19549 |
| C3 | 95721 | 9181 | 314 | 7229 | 16723 | 112444 | 16723 |
| C4 | 88300 | 150889 | 10239 | 49270 | 210398 | 219858 | 131558 |
@@ 8 Summary and Outlook
The modeling of signals from neutron interactions will benefit the DUNE low energy physics program in many ways. MeV-scale physics measurements using scintillation light in neutrino LArTPC are scarce. Efficient tagging of the neutron captures could improve energy resolution [APEX_lowE, MeVNueArCC]. Meanwhile, understanding the neutron capture signals in charge and light will help reject cavern neutron backgrounds. Developing such a neutron capture tagging technique at small LArTPC prototypes before DUNE FD operation is crucial for advancing the DUNE physics program [DUNE_lowE_1, DUNE_lowE_2, DUNE_lowE_3].

### v2

@@ 6.3 Discussion on Excess at High PE in Data
We further evaluate the PE spectra in three time periods, 0-160 s, 160-480 s, and 480-1050 s, in Fig. 15, Fig. 16, and Fig. 17, respectively. The first period roughly corresponds to the initial neutron beam-on period, where many neutron interactions, especially the inelastic interactions in active LAr, occurred. The second period (160-480 s) corresponds to time just after the neutron beam is off, where there are still many neutron activities. And the last period (480-1050 s) corresponds to the rest of the data acquisition window. In all cases, the high-PE excess is observed consistently during the whole DAQ window. Below, we discuss several possible sources that could produce the high-PE excess observed in the data in Fig. 10.
Cosmic background. The excess at high PE region could originate from residual cosmic flux not fully subtracted in PNS run data. To test this hypothesis, we plot the timing distribution separately for events with the number of PE below and above 1200 in Fig. 18. After subtracting the cosmic background in data, the not well modeled tail above 1200 PE accounts for less than 5% of the total neutron-induced light signals. For cosmic events, the timing distribution is expected to be flat as shown in Fig. 11. Instead, in the right plot of Fig. 18, for light signals with PE 1200, a decay time profile similar to the events with PE 1200 is observed (Fig. 18 left plot). This provides strong evidence that the excess at high PE is from neutron-related signals not modeled in the simulation.
@@
XA Absolute PDE. In the analysis (Sec. 5.2), a relative PDE calibration was performed to the 3% absolute PDE used in the simulation. However, the actual XA absolute PDE can vary by up to 50% when SiPMs are biased at higher overvoltages with a non-negligible cross-talk probability. The absolute PDE is also affected by the detailed XA configuration. A smaller absolute PDE in simulation would underestimate the number of PE and could partially explain the high PE excess in data. This is further discussed in Sec. 7 as a systematic effect. The precise measurement of absolute PDE for XA modules requires a dedicated laboratory setup and analysis. More details can be found in ref. [6].
@@ 7 Systematic Errors
Several important systematic effects are studied using the Fluka simulation. The uncertainty of the electric field in the buffer LAr region affects the scintillation photon yield. A simulation with zero electric field ("0 field") in this volume is evaluated, where the light yield reaches a maximum compared to the nominal electric field of 454 V/cm. The uncertainty of the initial production of scintillation light in buffer LAr and propagation to the active LAr readout volume is evaluated by setting the photon production in the buffer LAr volume to zero ("act only"). The uncertainty of the PNS position relative to the active volume inside the CB is modeled by shifting the PNS in the x (drift) and y (horizontal) directions by 5 cm ("dispx" and "dispy"). The uncertainty on the photon detector efficiency is studied by increasing it to 4.44% ("heff"), a value that corresponds to the product of efficiency and cross-talk probability for one of the device configurations described in ref. [6].
@@
The peak ADC amplitude to photoelectron calibration constant used in the data discussed in Sec. 5.1 affects the observed amount of light on each XA. The systematic shift of this constant up and down by one ADC unit is applied to the calculation in data, and the resulting counts are compared to the nominal case. The non-cosmic component in data has a max shift of 8.3% to the nominal scenario from the C4 XA module.
@@ 8 Summary and Outlook
The modeling of signals from neutron interactions will benefit the DUNE low energy physics program in many ways. MeV-scale physics measurements using scintillation light in neutrino LArTPC are scarce. Efficient tagging of the neutron captures could improve energy resolution [14, 8]. Meanwhile, understanding the neutron capture signals in charge and light will help reject cavern neutron backgrounds. Developing such a neutron capture tagging technique at small LArTPC prototypes before DUNE FD operation is crucial for advancing the DUNE physics program [7, 17, 13].